Using x^2=t,reduce the equation
2xydx+(1-x^2+2y)dy=0
to the linear form and hence solve it
1
Expert's answer
2020-10-07T17:46:13-0400

The given differential equation is

"2xy\ d x+(1-x^2+2y)dy=0" .............(1)

Let "x^2=t" .

On differentiation, "2x\ dx=dt"

Substituting these in Eq.(1) we get,

"ydt+(1-t+2y)dy=0" .This is the linear form of Eq.(1).

"ydt+(1+2y)dy-tdy=0\\
ydt-tdy+(1+2y)dy=0"

Dividing by "y^2" ,

"\frac{ydt-tdy}{y^2}+\frac{1+2y}{y^2}dy=0"

"d(t/y)+\left(\frac{1}{y^2}+\frac{2}{y}\right)dy=0" since "d(t/y)=\frac{ydt-tdy}{y^2}"

Integrating,

"\int d(t/y)+\int \left(\frac{1}{y^2}+\frac{2}{y}\right)dy=C" , where "C" is the constant of integration.

"\frac{t}{y}+\left(-\frac{1}{y} +2\ln y\right)=C"

Substituting "t=x^2,"

"\frac{x^2}{y}-\frac{1}{y}+2\ln y=C"

This is the solution of the differential equation (1).